# A Multiverse Outside of the Swampland

### v1

@@ 1 Introduction
The tension with inflation suggests another possible tension between the two conjectures and a construction that is relevant to some of the biggest open problems in particle physics: the Multiverse. Before illustrating the technical results of the paper, it is useful to recall a few characteristics of the Multiverse: 1) The physical processes that can create a Multiverse are simple and well-known [6, 7, 8, 9, 10], 2) The viability of this idea can be tested by searching for a UV completion of the SM with many metastable vacua and for a long period of inflation, 3) The Multiverse is the most concrete explanation that we have for the small value of the cosmological constant (CC), together with Weinberg’s argument [11].
Other than for the CC, the Multiverse can play an important role also in a number of other open questions in fundamental physics, where the most notable example is the value of the Higgs mass squared ($m_{h}^{2}$[12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
@@ 3 Eternal Inflation Outside of the Swampland
$-\frac{V^{\prime\prime}}{V}=\frac{M_{\rm Pl}^{2}}{f^{2}}\geq c^{\prime}\,.$ (10)
@@
We now have a simple model of eternal inflation compatible with our two working hypotheses in Eq.s (1) and (2). However, this model alone does not describe our Universe, as was noted for instance in [37]. What we imagine is that a second inflaton takes care of the last 60 $e$-folds of inflation, as discussed in Section 5. Here, we comment on why the simple model discussed in this Section is at odds with CMB measurements.
@@
To conclude this Section we comment on related work. First of all, eternal inflation can be excluded by a future measurement of positive curvature [38] that would exclude an epoch of eternal inflation followed by slow-roll. Planck is not sensitive enough to make a definitive statement [40].
Refer to captionRefer to caption
@@
Related analyses in the literature can be found in [37, 41, 42, 43]. In [37] the same model is considered and discarded because of the difficulties that we mentioned in explaining the CMB. In [41] they consider a linear potential near the top and do not tune, concluding that eternal hilltop inflation is incompatible with the refined dS conjecture. In [42] they consider eternal chaotic inflation and derive bounds on the $\mathcal{O}(1)$ parameters of the conjectures to accommodate it. It is possible only if, in Eq. (2), $c\lesssim 0.01$ and the field excursion is $\gtrsim 10^{2}M_{\rm Pl}$. Finally, in [43] an alternative way to have eternal inflation in a landscape without dS minima is discussed.
@@ 4 A Multiverse Without de Sitter Minima
We have seen that eternal inflation is consistent with both our working hypotheses Eq.s (1) and (2). But this is not enough to have a Multiverse that explains the CC and the weak scale, since we consider all minima in the landscape to be Minkowski or AdS, and their typical energy density to be $-M_{\rm Pl}^{4}$. In this landscape, natural initial conditions place us in a minimum with a CC of $\mathcal{O}\mathopen{}\mathclose{{}\left(-M_{\rm Pl}^{4}}\right)$. During the eternal inflationary period, we need to get to a minimum with higher vacuum energy in order to describe our Universe. Similarly, a typical minimum has $m^{2}_{h}\simeq M_{\rm Pl}^{2}$ and we need many (up)tunneling events to get to our Universe.
@@ 4.2 Dynamics of a Hawking-Moss Bubble
If we imagine that our Universes at higher vacuum energy are generated by a HM instanton, our first task is to compare the timescales for $\phi$ rolling down to its minimum and for the bubble wall evolution described in the previous Section. In Appendix B we find that the two are comparable. For simplicity, in the following we assume that the field rolls quickly to its minimum and then the bubble expansion starts. If the hierarchy is reversed, our qualitative arguments are not affected. In this latter case, to generate a baby Universe we have to impose the conditions on the mass of the bubble $M>M_{\textnormal{c}}$ not only on Hubble in the false vacuum $H_{\textnormal{F}}$, but also to Hubble at the top of the potential $H_{\rm top}$. This does not change our discussion as $H_{\rm top}>H_{\textnormal{F}}$ and this implies $M(H_{F}\to H_{\rm top})>M(H_{\textnormal{F}})$ as shown in Eq. (22). We can therefore impose our conditions on the critical mass using $H_{\textnormal{F}}$, which is relevant to the late-time evolution of the bubble, and they will be automatically satisfied also at early times.
@@ 5 The Last Sixty $\bm{e}$-Folds
In this Section we slowly build up towards two models that can give the last sixty $e$-folds of inflation without violating our working hypotheses. We find it instructive to start from the simplest possible models, point out general tensions with the Swampland conjectures, and then extend them while trying to minimize the model-building complexity. This constructive approach shows that it is not so easy to evade both working hypotheses at the same time, even from a purely bottom-up perspective. It also highlights generic tensions between slow-roll inflation and the two hypotheses, both in the single-filed and multifield cases.
@@ 5.2 Relaxing the Refined dS Conjecture
From Eq. 31, the number of $e$-folds is
@@ 5.4 Warm Inflation
The numerical coefficient $\kappa$ is $\mathcal{O}(100)$ and can be obtained on the lattice [79] as the parametric dependence of $Q$ on $T,\alpha$ and $f$, see also [80] for a review.
@@
The scalar power spectrum is dominated by temperature fluctuations [81] [^8] A different $Q$ dependence was obtained in [82], but it does not qualitatively change our results.
$\Delta_{R}^{2}\simeq\frac{15\sqrt{3}}{8\pi^{7/2}}\frac{\kappa\alpha^{5}H^{2}}{g_{*}f^{2}}\mathopen{}\mathclose{{}\left(\frac{Q(\xi)}{Q_{3}}}\right)^{9}\sqrt{Q(\xi)}\,,$ (49)
with $Q_{3}\simeq 7$. The tensor power spectrum is the usual one $\Delta_{h}^{2}\simeq\frac{2}{\pi^{2}}\frac{H^{2}}{M_{\rm Pl}^{2}}$, and $n_{s}-1$ in this model reads [78]
$n_{s}-1\simeq\frac{3}{7}\mathopen{}\mathclose{{}\left(27\epsilon_{Q}-19\eta_{Q}}\right)\,,$ (50)
@@
$f$ $\simeq M_{\rm Pl}\frac{Q_{3}^{3/2}(\Delta_{R}^{2})^{1/6}g_{*}^{2/3}\kappa^{1/2}\alpha^{5/2}}{Q(\xi)^{11/4}\xi}\,,$
$H$ $\simeq M_{\rm Pl}\frac{Q_{3}^{6}(\Delta_{R}^{2})^{2/3}g_{*}^{7/6}}{Q(\xi)^{15/2}\xi}\,.$
@@
$H\simeq 1.5\;{\rm GeV}\,,\quad f\simeq 10^{-8}\,M_{\rm Pl}\,,\quad\phi_{\textnormal{c}}\simeq\frac{2M_{\rm Pl}}{3}\,.$ (52)
@@
if we take $y$ perturbative and $M_{\rm UV}\lesssim M_{\rm Pl}$. This is just an example, but it illustrates a general problem of warm inflation that has to be circumvented by clever model-building as was done in the model [78] that we considered. There the coupling $\phi F\widetilde{F}$ can be written in terms of the derivative of a current $\phi\partial_{\mu}K^{\mu}$ and integration by parts shows that the $\phi$ potential does not receive any correction.
@@ Acknowledgments.
We thank A. Bedroya, M. Geller, L. Santoni, R. Wald for useful discussions and comments on the manuscript. This research is supported by a grant under the France And Chicago Collaborating in The Sciences (FACCTS) program. G. R. acknowledges funding from the European Unions Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie actions Grant Agreement no 945298-ParisRegionFP. The work of L.T. W. is also supported by DOE grant DE-SC-0013642.
@@ Appendix A Dynamics of False Vacuum Bubbles
The formal solution to this equation is well-known and takes the form
@@
This solution is not physical because the bubble forms in dS space and contains dS space, so we have horizons of size $\sim 1/H_{\textnormal{F,\,T}}$. For now we just note that Eq.s (84) and Eq. (85), albeit unphysical, are useful to identify the timescale of the bubble wall dynamics for an observer inside the bubble.

### v2

@@ 1 Introduction
The tension with inflation suggests another possible tension between the two conjectures and a construction that is relevant to some of the biggest open problems in particle physics: the Multiverse. Before illustrating the technical results of the paper, it is useful to recall a few characteristics of the Multiverse: 1) The physical processes that can create a Multiverse are simple and well known [6, 7, 8, 9, 10], 2) The viability of this idea can be tested by searching for a UV completion of the SM with many metastable vacua and for a long period of inflation, 3) The Multiverse is the most concrete explanation that we have for the small value of the cosmological constant (CC), together with Weinberg’s argument [11].
Other than for the CC, the Multiverse can play an important role also in a number of other open questions in fundamental physics, where the most notable example is the value of the Higgs mass squared ($m_{h}^{2}$[12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
@@ 3 Eternal Inflation Outside of the Swampland
$-\frac{M_{\rm Pl}^{2}V^{\prime\prime}}{V}=\frac{M_{\rm Pl}^{2}}{f^{2}}\geq c^{\prime}\,.$ (10)
@@
We now have a simple model of eternal inflation compatible with our two working hypotheses in Eqs. (1) and (2). However, this model alone does not describe our Universe, as was noted for instance in [38]. What we imagine is that a second inflaton takes care of the last 60 $e$-folds of inflation, as discussed in Section 5. Here, we comment on why the simple model discussed in this Section is at odds with CMB measurements.
@@
To conclude this Section we comment on related work. First of all, eternal inflation can be excluded by a future measurement of positive curvature [39] that would exclude an epoch of eternal inflation followed by slow-roll, or by a negative running of the spectral index [41, 42]. Planck is not sensitive enough to make a definitive statement [43].
Related analyses in the literature can be found in [38, 44, 45, 46, 47, 48, 49]. In [38] the same model is considered and discarded because of the difficulties that we mentioned in explaining the CMB. In [44] they consider a linear potential near the top and do not tune, concluding that eternal hilltop inflation is incompatible with the refined dS conjecture. In [45] they consider eternal chaotic inflation and derive bounds on the $\mathcal{O}(1)$ parameters of the conjectures to accommodate it. It is possible only if, in Eq. (2), $c\lesssim 0.01$ and the field excursion is $\gtrsim 10^{2}M_{\rm Pl}$. Finally, in [46] an alternative way to have eternal inflation in a landscape without dS minima is discussed.
@@ 4 A Multiverse Without de Sitter Minima
We have seen that eternal inflation is consistent with both our working hypotheses Eqs. (1) and (2). But this is not enough to have a Multiverse that explains the CC and the weak scale, since we consider all minima in the landscape to be Minkowski or AdS, and their typical energy density to be $-M_{\rm Pl}^{4}$. In this landscape, natural initial conditions place us in a minimum with a CC of $\mathcal{O}\mathopen{}\mathclose{{}\left(-M_{\rm Pl}^{4}}\right)$. During the eternal inflationary period, we need to get to a minimum with higher vacuum energy in order to describe our Universe. Similarly, a typical minimum has $m^{2}_{h}\simeq M_{\rm Pl}^{2}$ and we need many (up)tunneling events to get to our Universe.
@@ 4.2 Dynamics of a Hawking-Moss Bubble
If we imagine that our Universes at higher vacuum energy are generated by a HM instanton, our first task is to compare the timescales for $\phi$ rolling down to its minimum and for the bubble wall evolution described in the previous Section. In Appendix B we find that the two are comparable. For simplicity, in the following we assume that the field rolls quickly to its minimum and then the bubble expansion starts. If the hierarchy is reversed, our qualitative arguments are not affected. In this latter case, to generate a baby Universe we have to impose the conditions on the mass of the bubble $M>M_{\textnormal{c}}$ not only on Hubble in the false vacuum $H_{\textnormal{F}}$, but also to Hubble at the top of the potential $H_{\rm top}$. This does not change our discussion as $H_{\rm top}>H_{\textnormal{F}}$ and this implies $M(H_{\textnormal{F}}\to H_{\rm top})>M(H_{\textnormal{F}})$ as shown in Eq. (22). We can therefore impose our conditions on the critical mass using $H_{\textnormal{F}}$, which is relevant to the late-time evolution of the bubble, and they will be automatically satisfied also at early times.
@@ 5 The Last Sixty $\bm{e}$-Folds
In this Section we slowly build up towards two models that can give the last sixty $e$-folds of inflation without violating our working hypotheses. We find it instructive to start from the simplest possible models, point out general tensions with the Swampland conjectures, and then extend them while trying to minimize the model-building complexity. This constructive approach shows that it is not so easy to evade both working hypotheses at the same time, even from a purely bottom-up perspective. It also highlights generic tensions between slow-roll inflation and the two hypotheses, both in the single-field and multifield cases.
@@ 5.2 Relaxing the Refined dS Conjecture
From Eq. (31), the number of $e$-folds is
@@ 5.4 Warm Inflation
The numerical coefficient $\kappa$ is $\mathcal{O}(100)$ and can be obtained on the lattice [85] as the parametric dependence of $Q$ on $T,\alpha$ and $f$, see also [86] for a review. For an $\textnormal{SU}(N)$ confining gauge group, $\kappa$ is well approximated by $N^{5}$.
@@
In the strong regime the scalar power spectrum is dominated by temperature fluctuations [87, 88, 89]
$\Delta_{R}^{2}\simeq\frac{1}{4\pi^{2}}\frac{H^{3}T}{\dot{\phi}^{2}}\widetilde{F}_{\textnormal{M}}(Q)\,,$ (49)
where
$\widetilde{F}_{\textnormal{M}}(Q)$ $\equiv 5.4\times 10^{-5}Q^{7}$ (50)
[…]
The tensor power spectrum is the usual one $\Delta_{h}^{2}\simeq\frac{2}{\pi^{2}}\frac{H^{2}}{M_{\rm Pl}^{2}}$, and for large $Q$ $n_{s}-1$ in this model reads [89]
$n_{s}-1\simeq 8\epsilon_{Q}-6\eta_{Q}\,,$ (51)
@@
$f$ $\simeq M_{\rm Pl}\frac{(\Delta_{R}^{2})^{1/6}\kappa^{1/2}\alpha^{5/2}}{\widetilde{F}_{\textnormal{M}}(Q)^{1/6}g_{*}^{1/3}Q_{\textnormal{S}}^{7/6}}\,,$
$H$ $\simeq M_{\rm Pl}\frac{(\Delta_{R}^{2})^{2/3}g_{*}^{1/6}}{\widetilde{F}_{\textnormal{M}}(Q)^{2/3}Q_{\textnormal{S}}^{7/6}}\,.$
@@
$H\simeq 300\;{\rm GeV}\,,\quad f\simeq 10^{-8}\,M_{\rm Pl}\,,\quad\phi_{\textnormal{c}}\simeq\frac{5M_{\rm Pl}}{9}\,.$ (53)
@@
if we take $y$ perturbative and $M_{\rm UV}\lesssim M_{\rm Pl}$. This is just an example, but it illustrates a general problem of warm inflation that has to be circumvented by clever model-building as was done in the model [84] that we considered. There the coupling $\phi F\widetilde{F}$ can be written in terms of the derivative of a current $\phi\partial_{\mu}K^{\mu}$ and integration by parts shows that the $\phi$ potential does not receive any correction. For other attempts along these lines see for example [93, 94, 95]. For more on the relationship between warm inflation and the Swampland, see instead [96, 97, 98, 99].
@@ Acknowledgments.
We thank A. Bedroya, K. Berghaus, M. Geller, L. Santoni, R. Wald for useful discussions and comments on the manuscript. This research is supported by a grant under the France And Chicago Collaborating in The Sciences (FACCTS) program. G. R. acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie actions Grant Agreement no 945298-ParisRegionFP. The work of L.T. W. is also supported by DOE grant DE-SC-0013642.
@@ Appendix A Dynamics of False Vacuum Bubbles
The formal solution to this equation is well known and takes the form
@@
This solution is not physical because the bubble forms in dS space and contains dS space, so we have horizons of size $\sim 1/H_{\textnormal{F,\,T}}$. For now we just note that Eqs. (85) and Eq. (86), albeit unphysical, are useful to identify the timescale of the bubble wall dynamics for an observer inside the bubble.